### v1

# Edinburgh 2025/29

Abstract
We present a global determination of parton distribution functions (PDFs) that accounts for higher twist corrections in deep-inelastic scattering (DIS) and linear power corrections for single inclusive jet and dijet production data from the LHC. We determine these corrections and their associated correlated uncertainties using a methodology based on the theory covariance formalism, previously used to account for nuclear uncertainties and missing higher order uncertainties (MHOUs) in global PDF determinations. We then study the impact of the power corrections on the extracted PDFs, and demonstrate an improved description of the datam due to a reduced sensitivity to DIS data in the low-$x$ region where higher twist uncertainties are relatively large, and a reduced sensitivity to single inclusive jet data at relatively low $p_{T}$, where linear power corrections can be significant. Finally, we assess the impact of power corrections on observables relevant to LHC phenomenology, including Higgs production via gluon fusion, and the determination of $\alpha_{s}$. We find that these effects, while small, can be significant, improving perturbative convergence.
1 Introduction
The advent of the high energy frontier, now at a center-of-mass energy scale of 13.6 TeV at the LHC, offers an unprecedented opportunity to test the predictive power of the Standard Model. Remarkable progress has been made in calculations at next-to-next-to-leading order (NNLO) and even next-to-next-to-next-to-leading order (N3LO) accuracy in perturbative QCD, together with the inclusion of mass corrections, QED and Electroweak corrections. Nevertheless, some significant sources of theoretical uncertainty remain, both perturbative and non-perturbative, and in both the hard cross-sections themselves and in the parton distribution functions (PDFs). If our theoretical predictions are to match the accuracy and precision of LHC data, we need to control all theoretical uncertainties at the one percent level [Amoroso:2022eow].
A systematic estimation of all possible sources of significant theoretical uncertainty is important for two main reasons. First, it provides a more reliable assessment of the uncertainties associated with theoretical predictions, important for interpreting experimental data and establishing a robust framework for probing potential new physics. Second, and more subtly, since theoretical predictions at hadron colliders rely on accurate and precise PDFs, it is necessary to also account systematically for all the various theoretical uncertainties in the calculations used in global PDF fits.
Theoretical uncertainties arising from missing higher-orders (MHOUs) have been included in PDF fits using a theory covariance matrix [Ball:2018lag] constructed from scale variations [NNPDF:2019vjt, NNPDF:2019ubu, NNPDF:2024dpb], or by fitting nuisance parameters [McGowan:2022nag]. The covariance matrix approach has also been used to incorporate the non-perturbative uncertainties due to nuclear effects in global fits with deep-inelastic scattering (DIS) and Drell-Yan (DY) data on heavy nuclei and deuteron targets [Ball:2018twp, Ball:2020xqw].
Besides theoretical uncertainties due to perturbative corrections and nuclear effects, there are further uncertainties due to effects which are suppressed by a power of the hard scale. In DIS, the familiar expression for the cross-section, factorised into a convolution of a perturbative coefficient function and universal PDFs, is only the leading term in the operator-product expansion. Additional terms, known as higher twist (HT) corrections, are always present, but suppressed by powers of the hard scale $Q^{2}$. These HT corrections are in essence non-perturbative, and over the years there have been many attempts to determine them empirically by fitting $\mathcal{O}(\Lambda^{2}/Q^{2})$ corrections (here $\Lambda$ is the QCD scale) in the context of a PDF fit [Virchaux:1991jc, Alekhin:1998df, Alekhin:1999af, Martin:1998np, Martin:2003sk, Blumlein:2008kz, Alekhin:2012ig, Ball:2013gsa, Thorne:2014toa, Alekhin:2017kpj, Alekhin:2022tip, Cerutti:2025yji]. This is challenging, as the HT corrections can easily be confused with the unknown terms behaving as powers of $\mathcal{O}\left(1/\log(Q^{2}/\Lambda^{2})\right)$ arising due to MHOUs. Consequently most global PDF fits impose kinematic cuts on DIS data to ensure that the HT corrections are small.
For hadronic processes, such as DY or inclusive jets, there is no operator product expansion, but factorization still holds in the limit of large momentum transfer [Collins:1989gx]. Again there are non-perturbative contributions suppressed by powers of the hard scale $Q_{T}$. For sufficiently inclusive processes with a color singlet final state, such as DY, the power corrections (PCs) are expected to be $\mathcal{O}(\Lambda^{2}/Q_{T}^{2})$ [Beneke:1995pq, Caola:2021kzt], and thus small except close to threshold. However for many other hadronic observables there are generally linear power corrections of $\mathcal{O}(\Lambda/Q_{T})$. For example, linear PCs to jet processes have been studied extensively in $e^{+}e^{-}$ annihilation [Dasgupta:2003iq, Dasgupta:2007wa, Dasgupta:2008di, Cacciari:2007aq], with analytic estimates of the shift in the jet energy due to hadronization and underlying event. Studies of jets in hadronic collisions [Olness:2009qd] have extended these analyses to estimate the size of theoretical uncertainties related to various non-perturbative effects in inclusive single jet cross-sections. These estimates suggest that the linear PCs to jet processes might be significant even up to scales as large as $100$ GeV. Linear PCs are also expected in top production [Makarov:2023uet], unless one uses a short distance scheme for the top quark mass. Since linear PCs fall off much more slowly with hard scale than HT, it is even more difficult to disentangle them from MHOUs. To the best of our knowledge, an empirical determination of linear PCs to jets in the context of a global PDF fit has not yet been attempted.
In this paper, we use the theory covariance matrix formalism developed in [Ball:2018lag, Ball:2021icz], recently validated in Ref. [Ball:2025xgq], to provide reliable estimates of the corrections and theoretical uncertainties due to HT in DIS, and to PCs in single jet inclusive cross-sections and dijet cross-sections. The analysis is performed entirely within the open-source NNDPF framework [NNPDF:2021uiq]. The paper is organised as follows. In section 2 the methodology is discussed in detail. In section 3 we will present results, firstly for the HT, and then for PCs to inclusive jets and dijets. In section 4, we will include these corrections and uncertainties in a global PDF fit, assess its impact on the PDFs and relative luminosities, on the theoretical predictions for fundamental LHC processes such as the Higgs boson production, and on the determination of the strong coupling $\alpha_{s}$ from hadronic data. Finally, in section 5 we summarise our main conclusions and provide information on how to access the PDF fits with HT and PCs in LHAPDF format [Buckley:2014ana].
2 Methodology for the Determination of Power Corrections
In this section, we review the theory covariance methodology we will use to estimate the theoretical uncertainties associated with power corrections. This is based on the general formalism developed in Ref. [Ball:2021icz], where it is shown how to account for correlations between theory uncertainties in the PDFs and in the matrix element used in a prediction. When applied to parameters of physical interest, the methodology allows for the extraction of such parameters, together with their correlated uncertainties, in such a way as to account for all the correlations with the underlying PDFs. As such, the methodology was recently validated in a determination of the strong coupling [Ball:2025xgq], by comparing results to those obtained using the more cumbersome correlated replica method [Ball:2018iqk], and by closure testing [NNPDF:2017mvq]. In the following, we provide a brief overview of the theory covariance methodology before moving to its specific application to power corrections.
2.1 The Theory Covariance Method
The accurate and precise determination of PDFs requires careful consideration of the theoretical uncertainties associated with theoretical predictions. Theoretical uncertainties and their correlations can be incorporated in the fitting procedure through the implementation of a theoretical covariance matrix $C^{\textrm{th}}$. If we assume that these uncertainties are Gaussian and independent of the experimental uncertainties, the total covariance matrix $C$ can be written as the sum of the experimental $t_{0}$ covariance matrix [Ball:2009qv] $C^{\textrm{exp}}_{t_{0}}$ and a theory covariance matrix [Ball:2018lag, Ball:2018twp, Ball:2020xqw, NNPDF:2019vjt, NNPDF:2019ubu, NNPDF:2024dpb] incorporating in particular all MHOUs and nuclear uncertainties
$C_{ij}=C^{\textrm{exp}}_{t_{0},ij}+C^{\textrm{th}}_{ij}\,$, (2.1)
where the indices $i,j$ run over all the experimental data points.
In the NNPDF framework, the uncertainties encoded in this covariance matrix are propagated to the PDFs through a Monte Carlo approach. A set of $N_{\textrm{rep}}$ data replicas $D^{(k)}$ is generated satisfying
$\langle D_{i}^{(k)}\rangle=D_{i}\,,\qquad\langle(D_{i}^{(k)}-D_{i})(D_{j}^{(k)}-D_{j})\rangle=C_{ij}\,$, (2.2)
where $D_{i}$ denotes the central value of the experimental data, and the angle brackets denote an average over data replicas, in the limit $N_{\textrm{rep}}\to\infty$.
The metric $\chi^{2}$ used to optimise the set of PDF parameters $\theta$} is then
$\chi^{2}\left({\theta}^{(k)},D^{(k)}\right)=\frac{1}{N_{\textrm{dat}}}\sum_{i,j}^{N_{\textrm{dat}}}\left(T_{i}[{\theta}^{(k)}]-D_{i}^{(k)}\right)(C^{-1})_{ij}\left(T_{j}[{\theta}^{(k)}]-D_{j}^{(k)}\right)\,$, (2.3)
where $T_{i}$ is the theory prediction for data point $i$. The figure of merit in Eq. (2.3) ensures that the relative weights of the data points during the fit account for both experimental and theoretical uncertainties [Ball:2018lag]. The $\chi^{2}$ in Eq. (2.3) is then optimised for each replica $k$ in the Monte Carlo ensemble, using a cross-validation procedure to prevent overfitting. In terms of Bayesian reasoning, minimization of the $\chi^{2}$ corresponds to the maximization of the conditional probability of the predictions given the data
$P(T|D)\propto\exp\left(-\frac{1}{2}(T-D)^{T}C^{-1}(T-D)\right)\,$, (2.4)
where matrix notation has been adopted for simplicity.
Following Ref. [Ball:2021icz], additional sources of theoretical uncertainty can be modelled as fully correlated shifts in the predictions induced by a set of $N_{\textrm{nuis}}$ sources of theoretical uncertainty
$T({\lambda})=T+\sum_{\alpha=1}^{N_{\textrm{nuis}}}\lambda_{\alpha}\beta_{\alpha}\,$, (2.5)
where $\beta_{\alpha}$ are vectors (in the space of the data) of correlated shifts, and $\lambda_{\alpha}$ is the associated nuisance parameter that specifies the size of the shift. If we promote the nuisance parameters to stochastic variables, the conditional probability in Eq. (2.4) becomes
$P(T|D{\lambda})\propto\exp\left(-\frac{1}{2}(T+\sum_{\alpha=1}^{N_{\textrm{nuis}}}\lambda_{\alpha}\beta_{\alpha}-D)^{T}C^{-1}(T+\sum_{\alpha=1}^{N_{\textrm{nuis}}}\lambda_{\alpha}\beta_{\alpha}-D)\right)\,.$ (2.6)
Then by marginalising over the nuisance parameters, assuming for the latter a multivariate Gaussian distribution centred on zero with unit variance, the original expression in Eq. (2.4) is recovered, but with $C$ replaced by $C+S$, where
$S=\sum_{\alpha=1}^{N_{\textrm{nuis}}}\beta_{\alpha}\beta_{\alpha}^{T}\,$, (2.7)
is the theory covariance matrix for the $N_{\textrm{nuis}}$ additional sources of theoretical uncertainty. Hence, we can determine the optimal PDF with these new uncertainties by following the usual fitting procedure, but with $C$ replaced by $C+S$ in Eq.(2.2) and Eq.(2.3).
After determining this PDF, we can update the prior estimate of the nuisance parameters using information from the fit. Using Bayes’ theorem, the posterior distribution of the nuisance parameters is
$P({\lambda}|TD)\propto\exp\left(\frac{1}{2}\sum_{\alpha,\beta=1}^{N_{\textrm{nuis}}}(\lambda_{\alpha}-\bar{\lambda}_{\alpha})Z^{-1}_{\alpha\beta}(\lambda_{\beta}-\bar{\lambda}_{\beta})\right)\,$, (2.8)
which is a multivariate Gaussian distribution with mean
$\bar{\lambda}_{\alpha}(T,D)=-\beta_{\alpha}^{T}(C+S)^{-1}(T-D)\,$, (2.9)
and covariance
$Z_{\alpha\beta}=\delta_{\alpha\beta}-\beta_{\alpha}^{T}(C+S)^{-1}\beta_{\beta}\,.$ (2.10)
Note that Eq. (2.9) depends on the theoretical predictions, which in turn depend on the PDFs. In the context of PDF fits, theoretical predictions are thus statistical variables. In NNPDF the ensemble of PDF replicas is used as a proxy for the underlying distribution of PDFs, so the marginalization of the theoretical predictions over the nuisance parameters is performed by averaging the nuisance parameters over the PDF replicas:
$\bar{\lambda}_{\alpha}=\langle\bar{\lambda}_{\alpha}(T[\theta^{(k)},D)\rangle=-\beta_{\alpha}^{T}(C+S)^{-1}(T^{(0)}-D)\,$, (2.11)
where the central replica $T^{(0)}=\langle T[\theta^{(k)}]\rangle$. Likewise, the covariance matrix of the nuisance parameters is
$\bar{Z}_{\alpha\beta}=\delta_{\alpha\beta}-\beta_{\alpha}^{T}(C+S)^{-1}\beta_{\beta}+\beta_{\alpha}^{T}(C+S)^{-1}X(C+S)^{-1}\beta_{\beta}\,$, (2.12)
where $X$ is the covariance matrix of the PDF uncertainties of the theoretical predictions, defined as
$X=\langle(T^{(k)}-T^{(0)})(T^{(k)}-T^{(0)})^{T}\rangle\,.$ (2.13)
This contains the experimental uncertainties in the data, all the theoretical uncertainties of the theory covariance matrix used in the fit (including the new sources of uncertainty), and the uncertainty due to the methodology in the PDF fit.
From Eq. (2.11) we see that the nuisance parameters are informed by the fit, and in particular by the agreement between predictions and data weighted by the total covariance matrix. This is the information that the data provide on the nuisance parameters that is not absorbed by the PDFs. This is also evident from the expression for the covariance in the space of nuisance parameters, Eq. (2.12). This consists of three terms: the first is the prior, and the second gives the reduction in uncertainty due to the information from the data, assuming fixed PDFs. The third term then gives the increase in uncertainty due to the uncertainty in the PDFs. Note that
$0<Z_{\alpha\beta}\leq\bar{Z}_{\alpha\beta}\,$, (2.14)
since $(C+S)^{-1}X(C+S)^{-1}$ is positive semi-definite. Therefore, the nuisance parameters receive an extra uncertainty due to the PDF fit itself: there is less information on the theoretical uncertainties from the data because part of it is used in the fit to the PDFs. Note that Eq. (2.14) implies that when the PDF uncertainty $X$ is large, then the covariance of the nuisance parameters can become larger than the prior.
We now show how the formalism above can be adapted for the present study. We model the shifts to predictions due to power corrections through a functional form that depends on a set of $N_{h}$ parameters $h_{\alpha}$. The details of the functional form depend on the physical process under consideration, and will be discussed in the next subsection. What is important here is that the power correction induces a shift in the theoretical predictions which is linear in $h_{\alpha}$:
$T(\{h\})=T+\sum_{\alpha}h_{\alpha}T^{\prime}_{\alpha}\equiv T+\sum_{\alpha}\lambda_{\alpha}\beta_{\alpha}\,$, (2.15)
so if we choose for each $\alpha$
$\beta_{\alpha}=\delta h_{\alpha}T^{\prime}_{\alpha}\,,\qquad h_{\alpha}={\lambda_{\alpha}}{\delta h_{\alpha}}\,.$ (2.16)
and take the prior distribution of the nuisance parameters as unit Gaussians centred on zero, as before, then $h_{\alpha}$ are also Gaussian variables centred on zero, but with prior covariance $H^{\rm prior}_{\alpha\beta}=(\delta h_{\alpha})^{2}\delta_{\alpha\beta}$.
The theory covariance matrix $S$ to be used in the fit will then be constructed as in Eq. (2.7) using the shifts in Eq. (2.16). Note that since the corrections in Eq. (2.15) are linear in the nuisance parameters, there is no need (unlike in Refs. [AbdulKhalek:2019bux, AbdulKhalek:2019ihb, NNPDF:2024dpb, Ball:2025xgq]) to consider contributions from positive and negative shifts separately, nor to consider complicated prescriptions: Eq. (2.7) is perfectly sufficient. After performing the fit including the new theory covariance matrix, we can compute the posterior for the nuisance parameters using Eq. (2.11), and their covariance using Eq. (2.12), which after the rescaling (2.16) gives the posterior parameters $h^{\rm post}_{\alpha}$ and their covariance
${h}^{\rm post}_{\alpha}=\delta h_{\alpha}\bar{\lambda}_{\alpha}\,,\qquad\textrm{H}^{\rm post}_{\alpha\beta}=\delta h_{\alpha}\delta h_{\beta}\bar{Z}_{\alpha\beta}\,$, (2.17)
and thus the power correction. This technique can be applied equally to determine HT corrections to DIS and power corrections to hadronic observables, as we now discuss.
2.2 Modelling Higher Twist Corrections in DIS
Power corrections in predictions for DIS observables arise from higher twist terms (twist four and above) in the operator product expansion, which are suppressed by powers of $Q^{2}$ compared to the leading twist (twist two). Although HT terms can be factorised into perturbative coefficients functions and non-perturbative generalised parton distribution functions, the latter are at present unknown. The HT terms are thus a source of systematic uncertainty that affects all DIS theoretical predictions, becoming particularly relevant at low $Q^{2}$ and large $x$ (low invariant mass of the final state $W^{2}$). Here we will only consider twist four terms, of $O(\Lambda^{2}/Q^{2})$. Where these are small, it is reasonable to assume that twist six terms of $O(\Lambda^{4}/Q^{4})$ will be negligible.
The DIS data we consider are structure function data taken on proton $F_{2}^{p}$ and deuteron $F_{2}^{d}$ targets, ratio data $F_{2}^{d}/F_{2}^{p}$, reduced cross-sections on proton targets, both neutral current (NC) and charged current (CC), and CC reduced cross-section on nuclear targets. We do not include DIS data on charm or bottom production in this study, nor data on DIS+jets.
There are two distinct approaches to modelling higher twist corrections to theoretical predictions for DIS data: the additive approach [Alekhin:1998df, Alekhin:1999af, Blumlein:2008kz, Alekhin:2012ig, Alekhin:2017kpj, Alekhin:2022tip], and the multiplicative approach [Virchaux:1991jc, Martin:1998np, Martin:2003sk, Ball:2013gsa, Thorne:2014toa]. In the additive approach one ignores all perturbative corrections to the HT. In the multiplicative approach, one first factors out the leading twist contribution, so in effect one assumes that the HT evolves perturbatively in the same way as the leading twist. Both approaches are thus approximate, and which one adopts is largely a matter of taste (see Ref. [Cerutti:2025yji] for a recent phenomenological comparison).
We consider HT corrections as multiplicative shifts to DIS observables, which comprise measurements of structure functions $F_{2}$ and reduced cross-sections. We model the shifts as functions of Bjorken-$x$, to be determined empirically. Moreover, given that in principle HT can depend on nuclear effects, the shifts for proton and deuteron targets are parametrised independently [Cerutti:2025yji]. Hence, we write the measured structure functions as
$F^{p}_{2}(x,Q^{2})$ $=F^{p,\textrm{LT}}_{2}(x,Q^{2})\left(1+\frac{H_{2}^{p}(x)}{Q^{2}}\right)\,$, (2.18)
$F^{d}_{2}(x,Q^{2})$ $=F^{d,\textrm{LT}}_{2}(x,Q^{2})\left(1+\frac{H_{2}^{d}(x)}{Q^{2}}\right)\,.$ (2.19)
Since the corrections are multiplicative, we can correct the HERA NC reduced cross-sections using the same factor:
$\sigma_{\rm NC}(x,Q^{2})=\sigma^{\textrm{LT}}_{\rm NC}(x,Q^{2})\left(1+\frac{H_{2}^{p}(x)}{Q^{2}}\right)\,.$ (2.20)
For ratio data, we have
$\frac{F_{2}^{d}}{F_{2}^{p}}=\left(\frac{F_{2}^{d}}{F_{2}^{p}}\right)^{LT}\left(1+\frac{H_{2}^{(d)}-H_{2}^{(p)}}{Q^{2}+H_{2}^{(p)}}\right)\,.$ (2.21)
We have checked that the denominator in the HT correction can be replaced with $Q^{2}$, consistent with our assumption that twist six terms (terms of $O(\Lambda^{4}/Q^{4})$) can be ignored.
Finally, DIS CC cross-sections are corrected similarly, by introducing a common multiplicative shift for CHORUS, NuTeV, and HERA CC data regardless of the nuclear target used in the experiment. Thus we will assume that nuclear effects at higher twist in CC data are similar to those included at leading twist [Ball:2018twp]. This approximation will be justified a priori. Hence, the shift for any CC cross-section will be written as
$\sigma_{\rm CC}(x,Q^{2})=\sigma^{\textrm{LT}}_{\rm CC}(x,Q^{2})\left(1+\frac{H_{\rm CC}(x)}{Q^{2}}\right)\,.$ (2.22)
The various functions $H(x)$ that model the shifts are parametrised as piecewise linear functions, interpolating between a grid of points in $x$. For the present analysis, points are placed at
$x\in[0.0,0.001,0.01,0.1,0.3,0.5,0.7,0.9,1]$
for all structure functions and targets. We further assume that $H(1)=0$, but make no assumption about $H(0)$. We thus write
$H(x)=\sum_{\alpha=1}^{8}h_{\alpha}f_{\alpha}(x)\,$, (2.23)
where the $f_{\alpha}(x)$ are piecewise linear interpolating functions. The advantages of having a linear dependence on the parameters $h_{\alpha}$ is twofold: it has a simple physical interpretation of the parameters used in the interpolation, and it simplifies the error propagation since Eq. (2.15) becomes an exact result. The vectors $T_{\alpha}^{\prime}$ are then simply proportional to the vectors $T$, that is $T_{\alpha}^{\prime}=Tf_{\alpha}(x)/Q^{2}$.
Finally, it is important to note that we consider the shifts to be fully correlated amongst all the observables that share the same HT correction. So, for example, a shift to $F_{2}^{p}$ will affect all the observables that depend on $F_{2}^{p}$, whether they are pure structure functions, the NMC ratio, or the NC reduced cross-sections.
2.3 Modelling Linear Power Corrections for Jet and Dijet Cross-sections
In NNPDF4.0 we include data on both single-inclusive jet and dijet cross-sections, from ATLAS and CMS. These are compared to factorised perturbative calculations of the form (for a single-inclusive jet cross-section)
$\left[\frac{d\sigma_{\textrm{jet}}}{dp_{T}dy}\right]_{\text{pert}}=\sum_{p,q}\int dx_{1}dx_{2}f_{p}(x_{1})f_{q}(x_{2})\frac{d\hat{\sigma}_{pq\rightarrow\textrm{jet}}}{dp_{T}dy}.$ (2.24)
Here $f_{p}(x_{1})$ and $f_{q}(x_{2})$ are the PDFs of the incoming partons $p$ and $q$, and $d\hat{\sigma}_{pq\rightarrow\textrm{jet}}/dp_{T}dy$ is the partonic cross-section computed at NNLO in perturbative QCD. MHOUs estimated from scale variations have been accounted for in a variant of NNPDF4.0 fits [NNPDF:2024dpb]. While it has been understood for some time that there are linear power corrections to these processes, from a wide variety of sources, in global PDF fits these are generally ignored.
There are many factors that influence the measurement of jet cross-sections, mainly to do with the way the jets and their associated energy scales are defined. For instance, the algorithm used to identify the jets introduces a dependence on the radius of the jet $R$. Furthermore, assuming that jets have been faithfully reconstructed, there is a variety of physical phenomena, not all belonging to the domain of perturbative calculations, that can affect the measured energy of the jet. In other words, the energy of the jet does not necessarily match the energy of the hard parton that initiated the jet. For example, the underlying event originates from uncontrolled interactions in the initial proton-proton collision (multiple parton interactions and colour interactions), as well as contributions from initial- and final-state radiation. As a results, the measured energy of the jet is shifted due to the external contribution from the underlying event that flows into the jet. Investigations to disentangle the underlying event contribution from single-inclusive jets exploiting Monte Carlo event generators have been performed by ATLAS [ATLAS:2014iez] and CMS [CMS:2013ycn]. On the other hand, in the process of hadronization, some of the jet energy leaks out from the jet cone, reducing the energy that is eventually measured. This process is modelled through the use of final state Monte Carlo generators. Experimentalists correct for these effects in their measurements, with associated systematic uncertainties.
However in addition to all this, the theoretical calculation of the perturbative jet cross-section is also subject to non-perturbative uncertainties, resulting in power corrections of $O(\Lambda/p_{T})$ to the parton level transverse momentum and thus to the perturbative cross-sections. These genuine non-perturbative effects are by their nature very difficult to disentangle from the quasi-partonic effects already corrected for in the experimental measurements and thus, as with higher twist, an empirical approach is the only practical way to assess them.
Whatever the source of the effect, a change in energy results in a shift $\delta p_{T}$ to the fixed parton level transverse momentum $p_{T}$ of the jet. The correction $\delta p_{T}$ has been estimated in Refs. [Dasgupta:2007wa, Dasgupta:2008di, Cacciari:2007aq]. Since jet cross-sections fall very steeply with $p_{T}$ (typically as $\sim p_{T}^{-n}$, with $n$ as large as six or seven), such a shift results in a multiplicative power correction of the form $1+n\delta p_{T}/p_{T}$, which may be large simply because $n$ is large [Olness:2009qd].
We will thus model the shift in the single-inclusive jet cross-section induced by power corrections as
$\frac{d\sigma_{\textrm{jet}}}{dp_{T}dy}=\left[\frac{d\sigma_{\textrm{jet}}}{dp_{T}dy}\right]_{\text{pert}}\biggl(1+\frac{H_{j}(y)}{p_{T}}\biggr)\,$, (2.25)
where the perturbative contribution is computed in perturbative QCD using the NNLO factorization (2.24) and the associated MHOU. We allow the function $H(y)$ to depend on the jet rapidity $y$, but not on the hard scale $p_{T}$, just as in the DIS case where the HT functions depend on $x$ but not $Q^{2}$. Note that for jets a multiplicative parametrization is mandatory, since, unlike for DIS, the perturbative cross-section is falling steeply as the hard scale increases.
Similar arguments apply to dijet cross-sections, where the hard scale is now the invariant mass $m_{jj}$ of the jets. The power corrections for the differential cross-section for dijets can then be modelled as
$\frac{d\sigma_{\textrm{2jet}}}{d\eta dm_{jj}}=\left[\frac{d\sigma_{\textrm{2jet}}}{d\eta dm_{jj}}\right]_{\text{pert}}\biggl(1+\frac{H_{jj}(\eta)}{m_{jj}}\biggr)\,$, (2.26)
where $\eta$ is either the rapidity separation $y^{*}=|y_{1}-y_{2}|/2$ (in ATLAS) or the maximum absolute rapidity $y_{\textrm{max}}=\rm{max}(|y_{1}|,|y_{2}|)$ in CMS.
The functions $H_{j}(y)$, $H_{jj}(\eta)$ that model the shifts are parametrised as piecewise linear functions just as for DIS (see Eq. (2.23)), though here we only use three points in $\eta$,
$\eta\in[0.0,1.5,3.0]$
since the rapidity dependence is expected to be rather weak. We remark that while for single-inclusive jets we use the same function $H_{j}(y)$ for both ATLAS and CMS, for dijets we must use different functions, since the data are analysed using different rapidity variables.
Note that since the linear power corrections fall off much more slowly with the hard scale than the HT corrections to DIS, they will be even more difficult to disentangle from MHO perturbative corrections. This problem is compounded by the fact that the N3LO corrections to jet processes are unknown, to the perturbative predictions are less precise than for DIS. On the other hand, since as explained in [Olness:2009qd] the scale of the linear power corrections may be relatively large (several GeV rather than several hundred MeV), they may continue to be significant up to relatively high scales. They can thus be significant, even at the LHC.
3 Determination of Higher Twist and Jet Power Corrections
In this section, we first discuss the determination of the HT corrections to DIS, and then apply the same technique to the determination of PCs for single inclusive jets and dijets. All these results will be obtained using the theory covariance methodology described in the previous section.
The fits presented in this section and the next have been produced using the open-source NNPDF framework [NNPDF:2021uiq] based on the same dataset and theory parameters as used in NNPDF4.0 [NNPDF:2021njg] with MHOUs as in Ref. [NNPDF:2024dpb] and aN3LO corrections as in Ref. [NNPDF:2024nan]. Theory predictions have been generated with the pipeline of Ref. [Barontini:2023vmr], which supersedes the tools used for NNPDF4.0. In particular, interpolation grids are generated using PineAPPL [Carrazza:2020gss] with the DGLAP evolution computed using EKO [Candido:2022tld] and DIS observables using the Yadism package [Candido:2024rkr]. For a detailed discusion of the updated theory pipeline see App A of Ref. [NNPDF:2024djq].
3.1 Determination of Higher Twist
The DIS measurements included in NNPDF4.0 [NNPDF:2021njg] consist of fixed-target NC structure function data from SLAC [Whitlow:1991uw], BCDMS [Benvenuti:1989rh], and NMC [Arneodo:1996kd, Arneodo:1996qe], fixed-target inclusive and dimuon CC cross-sections from CHORUS [Onengut:2005kv] and NuTeV [Goncharov:2001qe, MasonPhD], and collider NC and CC cross-sections from HERA [Abramowicz:2015mha]. In NNPDF4.0 (and indeed all earlier NNPDF global fits) cuts are applied to all DIS data, $Q^{2}>3.49~{\rm GeV}^{2}$, $W^{2}>12.5~{\rm GeV}^{2}$, in order to avoid the kinematic regions where the HT is expected to be significant.
All calculations of DIS processes in NNPDF4.0 include target mass corrections [Schienbein:2007gr], following the implementation of Refs. [Georgi:1976ve, Ball:2008by], heavy quark mass corrections implemented using FONLL [Forte:2010ta, Ball:2011mu, Barontini:2024xgu], fitted charm [Ball:2015tna, Ball:2015dpa, Ball:2016neh], and nuclear corrections and uncertainties, both for heavy nuclei [Ball:2018twp] and deuteron targets [Ball:2020xqw].
We estimate the higher twist using the theory covariance methodology, as described in section 2, constructing a prior theory covariance matrix by computing the shifts in DIS predictions for a prior variation of the HT parameters $h_{\alpha}$. Since our prior knowledge of the HT is limited, we assume that the prior is centred on zero and allow the higher twist nodes $h_{\alpha}$ to vary independently in the range $\pm 1~\rm{GeV}^{2}/(1-x)$. This choice is motivated by the expectation that HT corrections are enhanced at large-$x$ due to the fact that the hadronic mass of the final system approaches threshold, which amplifies non-perturbative effects in this region. The prior distributions of the three HT functions $H_{2}^{p}(x)$, $H_{2}^{d}(x)$ and $H_{CC}(x)$ are shown in Fig. 3.1.
Figure 3.1: The three HT functions $H_{2}^{p}(x)$ (top), $H_{2}^{d}(x)$ (middle) and $H_{CC}(x)$ (lower). The prior uncertainty is shown in gray. The posterior distribution is then shown for HT determinations where the leading twist is computed at NNLO (blue), NNLO+MHOU (green) and aN3LO+MHOU (red).
We then add this prior theory covariance matrix to the covariance matrix $C_{ij}$ used in NNPDF4.0, perform a new fit, and use the results of this fit to extract posterior HT parameters. We then check that when the prior is varied (by a factor of two in either direction), the posterior HT remains unchanged. To maximise the sensitivity to HT corrections, we perform the fit in an extended kinematic region that includes DIS data that we expect to be sensitive to the non-perturbative effects that we are trying to capture. For that reason, we lower the standard cuts to $Q^{2}>2.5$ GeV2 and $W^{2}>3.24$ GeV2. The impact of this lowered cut on the DIS dataset is shown in Fig. 3.2.
Figure 3.2: The $x,Q^{2}$ plane of the DIS data points used in the fits: the NC data (left) and CC data (right). The solid line indicates the position of the standard cut used in NNPDF4.0. The additional data points below this line are the data added when the cuts are lowered to determine the HT. The size of the points indicates the size of the the absolute value of the relative shift due to HT, determined in the NNLO fit. The colour code represents the ratio between the posterior uncertainty due to HT and the experimental uncertainty.
The HT determined in this way is shown in Fig. 3.1 for three different sets of PDFs: NNLO, NNLO with MHOUs, and aN3LO with MHOUs. At NNLO we find that the posterior distribution is significantly more constrained than the prior, justifying the choice of prior. For the proton structure functions, the HT correction $H_{2}^{p}$ is constrained to be very small at low and intermediate $x$, consistent with zero within uncertainties, but increases towards larger values of $x$, reaching a peak of about $4$ GeV2. The deuteron HT correction $H_{2}^{d}$ shows a similar pattern, although with slightly larger uncertainties. Note that, in both cases, the low-$x$ region is well constrained by the data. The HT for CC observables, $H_{CC}$, shows a similar pattern, but is less well constrained, particularly at very small $x$ and very large $x$, where there is no CC data. In particular, at $x=0.9$ there is no constraint at all, with the posterior equal to the prior. This said, we see no evidence here for the expectation [Dasgupta:1996hh], based on renormalons, that HT effects for CC processes are larger than for NC processes, or that they have a different shape.
In Fig. 3.3 we show a comparison of the HT for the $F_{2}$ structure function in proton and deuteron determined in our global NNLO fit, with results obtained in an early PDF fit to just SLAC and BCDMS data by Virchaux and Milsztajn [Virchaux:1991xp]. While this fit was only at NLO, it used a similar multiplicative parametrization to our own, and target mass corrections were included, so the comparison is meaningful. Considering the fact that this fit was performed before the advent of NMC and HERA data, and all the technical improvements made over the intermediate thirty years, the agreement is rather impressive, though it should not come as too much of a surprise since it is the same data (SLAC and BCDMS) that dominates the large-$x$, low-$Q^{2}$ region today. Direct comparisons with more recent determinations which use an additive HT parametrization [Alekhin:1998df, Alekhin:1999af, Blumlein:2008kz, Alekhin:2012ig, Alekhin:2017kpj, Alekhin:2022tip] are more difficult, since the results necessarily depend to some extent on the choice of scale. However for these too, the results for HT correction to $F_{2}^{p}$ are qualitatively consistent with our own, within uncertainties.
Figure 3.3: Comparison of our NNLO $F_{2}$ higher twist to that determined in protons and deuteron by Virchaux and Milsztajn in a NLO PDF fit to SLAC and BCDMS data [Virchaux:1991xp]. They include TMCs, and use the same multiplicative parametrization of the HT, fitting proton and deuteron data separately, so the results are directly comparable to our own.
When we add MHOUs to our NNLO fit, the HT correction is generally reduced in size. The same happens again when we go to aN3LO, where the MHOUs are significantly reduced. This shows that some of the supposed HT being seen at NNLO is in fact due to MHOUs, principally the N3LO perturbative correction to the leading twist contribution. This shows the difficulty in distinguishing phenomenologically between $Q^{2}$ dependence which is power like, and $Q^{2}$ dependence which is merely a power of a logarithm. This effect was also seen when going from NLO to NNLO in earlier attempts to determine HT [Martin:2003sk]. One is forced to conclude that even the HT determined in the aN3LO fit may still be contaminated by a perturbative contribution. Similarly, $H_{2}^{d}$ may be tangled with the deuteron nuclear correction [Ball:2020xqw], and $H_{CC}$ with the nuclear corrections applied to much of the CC data [Ball:2018twp].
The size and significance of the HT we determine is also displayed in the kinematic plots in Fig. 3.2, where the per-datapoint impact of the posterior distribution for NC and CC datasets is shown (indicated by the size of each data point), and its uncertainty compared to the overall experimental uncertainty (indicated by the colour of each data point). For both NC and CC corrections, we see that the relative shift of the HT exhibits the expected behaviour, being largest at low $Q^{2}$ values where the non-perturbative effects dominate, and also increasing in size at large $x$. Most importantly, the size of the relative shifts increases significantly for both NC and CC DIS datasets, particularly in the cut region, up to about  10%. This justifies the cuts chosen in NNPDF4.0, designed precisely to eliminate data significantly contaminated with HT contributions. For the large-$x$ data that remain above the cut, the effect of the HT is small, and in particular the uncertainty due to the HT is rather smaller than the (quite large) experimental uncertainties. However, for the NC data from HERA at small $x$, although the HT is small, the uncertainty due to the HT is a significant fraction of the experimental uncertainty, even for some data above the cut. This is because the HERA experimental uncertainties are themselves small. This effect is reduced when MHOUs are added at NNLO, and further reduced at aN3LO. This last observation should be borne in mind when considering the effect of HT on the PDFs in the next section.
3.2 Determination of Linear Power Corrections to Jets and Dijets
We now turn to the determination of power corrections for jet observables. The jet cross-sections included in the NNPDF4.0 fit are double differential single-inclusive jet data at $\sqrt{s}=8$ TeV and dijet cross-sections at $\sqrt{s}=7$ TeV from ATLAS [ATLAS:2017kux, ATLAS:2013jmu] and CMS [CMS:2016lna, CMS:2012ftr]. No kinematic cuts were applied to these data in NNPDF4.0. An internal cut on the jet $p_{T}$ was applied to the single-jet measurements of CMS in order to avoid points for which $p_{T}\lesssim 70$ GeV, where fixed-order calculations are not reliable. This cut is not necessary for ATLAS single-jet data since the lowest $p_{T}$ bin starts at $70$ GeV. Finally, no additional cuts are applied to dijet data, as the lowest value of the hard scale is one order of magnitude larger than that for the single-inclusive jet data. These considerations will become relevant in the discussion below.
As discussed in Sec. 2, we use the same methodology as for the higher twist determination, but with a different prior. For jet observables, the prior is still centred on zero, but given the expectation that $H(\eta)$ might be relatively large, we allow the shift to vary in the range [20.0,20.0][-20.0,20.0] GeV. As for the HT, we check that the results are independent of the prior by varying it by a factor of two in either direction.
In Fig. 3.4 we present the posterior distribution for the power corrections to single inclusive jets and dijets, and in Fig. 3.5 the related kinematic distribution of the data points. We observe two distinct patterns for single-inclusive jets and dijets. Starting with the single-inclusive jets, we find that not only is the prior significantly constrained, but also that the posterior distribution is peaked around $10$ GeV. Further, the posterior shows a weak dependence on the rapidity. It should be noted that the previous study in Ref. [Olness:2009qd] estimated a positive power correction of about $7$ GeV for single-inclusive jet cross-sections, independent of rapidity. Our result is consistent with this estimate. For the dijets, we find only weak evidence for a linear PC in the NNLO fit, although uncertainties remain large, showing that the data provide only weak constraints on the PCs. Note that for the dijet kinematics the bins with lowest invariant mass (and thus most likely to be affected by PCs) are at central rapidity, and even here the lowest invariant mass bin is still at a rather higher scale than the lowest $p_{T}$ scale of the single inclusive jet data.
Figure 3.4: Same as Fig. 3.1, but for single-inclusive jets (top), ATLAS dijet (lower left) and CMS dijet (lower right) observables.
Figure 3.5: The $y,p_{T}$ plane of the single-inclusive jet data points (left) and the $y,m_{jj}$ plane of the dijet data points (right). For the dijet data, the ATLAS and CMS definitions of rapidity are both used on the same plot. As in Fig. 3.5, the size of the points indicates the size of the the absolute value of the relative shift due to the PCs, determined in the NNLO fit. The colour code represents the ratio between the posterior uncertainty due to PCs and the experimental uncertainty.
There is perhaps a small reduction in the size of the PC to single-inclusive jets when MHOUs are added at NNLO, particularly at large rapidities, but in the aN3LO fit the result is entirely consistent with NNLO. This is not too surprising, since the aN3LO fit does not include an N3LO correction to the perturbative cross-section, since such corrections are at present unknown. However for the dijets, adding MHOUs or going to an aN3LO fit seems to give a more negative power correction at central rapidities (for both ATLAS and CMS), even though the uncertainties remain large. It is possible that this large negative PC at central rapidity will be cancelled, at least in part, by the (as yet unknown) N3LO correction to the perturbative cross-section.
The size of the shifts due to the linear PC are shown in Fig. 3.5 by the size of the data points. Clearly the power corrections to jets fall off slowly with $p_{T}$, as expected from the $1/p_{T}$ scaling behaviour, but can be as large as about $14\%$ for the lowest $p_{T}$ values (around $75$ GeV) of the single-inclusive jet data, across all rapidities. The uncertainty due to the linear PC is quite a modest fraction of the experimental uncertainty, even at these low scales, since the experimental uncertainty is already quite large. The dijet data start at invariant dijet masses of around $240$ GeV, and this only for the bins at central rapidity: for the large rapidity bins the invariant mass is always above a TeV. This explains why at the LHC dijet data are less vulnerable to power corrections than single-inclusive jet data. The underlying reason for this is kinematic: it is easier to identify unambiguously the hard scale for dijets than for single-inclusive jets.
4 The NNPDF4.0HT Parton Sets
We evaluate the phenomenological consequences of including the HT corrections and PCs discussed in the previous sections at the level of a new global PDF fit, and discuss the impact on relevant LHC observables.
4.1 Global PDF fits with Power Corrections
The HT and power corrections discussed above were determined by performing global PDF fits to the NNPDF4.0 dataset, but lowering the kinematic cuts to increase sensitivity to the HT. However, we do not suggest performing a fit with these lowered cuts to determine the PDFs, since such a fit would include data which is significantly affected by HT. Instead, we use the HT and the PCs determined above to give corrections to the theory predictions in a fit with the conventional cuts used in NNPDF4.0 [NNPDF:2021njg], i.e. $Q^{2}>3.49$ GeV2 and $W^{2}>12.5$ GeV2.
The multiplicative factors are constructed following the definition of the shifts in eqs. 2.19, 2.20, 2.21 and 2.22 for DIS and in eqs. 2.25 and 2.26 for jet observables, using the central values of the posterior parameters, $h^{\rm post}_{\alpha}$, eq. 2.17. Thus the shift in theoretical predictions is given by
$\delta T=\sum_{\alpha}\bar{\lambda}_{\alpha}\beta_{\alpha}.$ (4.1)
Likewise the uncertainties and correlations of these parameters are encoded in the posterior covariance matrix $H^{\rm post}_{\alpha\beta}$, eq. 2.17, which propagates through eq. 2.15 into a covariance matrix in the space of data which gives the uncertainties and correlations due to the HT and PCs:
$C^{\rm post}=2\sum_{\alpha,\beta}\beta_{\alpha}\beta_{\beta}^{T}\bar{Z}_{\alpha\beta}.$ (4.2)
Here we have introduced a factor of two to provide a conservative estimate of the posterior uncertainty: we do not want to be too aggressive with the uncertainty on the HT and PCs, in part because some data have been used twice (first to determine the PCs, then again to determine the PDFs), but also because of the inherent difficulties discussed previously in finding clear evidence for HT and PCs. The total covariance matrix $C$ to be used in the PDF fit is then obtained by adding this contribution to the existing covariance matrix eq. 2.1, which already includes the experimental uncertainties, and other sources of theoretical uncertainty (nuclear and MHO). We use this setup to perform a global PDF fit to the NNPDF4.0 dataset for each perturbative theory setting considered in this work. For each setting, the multiplicative shifts and the posterior uncertainty determined in section 3 are chosen consistently with the theory setting employed in the PDF fit. We discuss the results in the following sections.
4.2 Fit Quality
Dataset $_{\rm data} NN$LO NNLO+MHOU aN3LO+MHOU
baseline only shift shift+unc. baseline only shift shift+unc. baseline only shift shift+unc.
DIS NC 2100 1.230 1.228 1.211 1.204 1.200 1.195 1.203 1.201 1.196
DIS CC 989 0.902 0.912 0.900 0.901 0.898 0.897 0.916 0.920 0.912
DY NC 735 1.201 1.184 1.184 1.150 1.156 1.150 1.159 1.140 1.143
DY CC 157 1.470 1.486 1.460 1.366 1.347 1.317 1.370 1.405 1.383
Top pairs 64 1.250 1.32 1.293 1.440 1.459 1.381 1.407 1.347 1.310
Single-inclusive jets 356 0.954 0.928 0.942 0.809 0.787 0.791 0.832 0.808 0.804
Dijets 144 2.018 1.928 1.929 1.713 1.479 1.485 1.674 1.480 1.477
Prompt photons 53 0.756 0.735 0.734 0.670 0.686 0.688 0.680 0.696 0.691
Single top 17 0.364 0.372 0.374 0.376 0.384 0.380 0.355 0.354 0.353
Total 4615 1.171 1.164 1.154 1.135 1.123 1.116 1.140 1.130 1.124
Table 4.1: The $\chi^{2}$ per data point for fits to the global NNPDF4.0 dataset with conventional cuts, with and without HT and PCs, and with the addition of the HT and PC uncertainty. The total is broken down into the contributions from the various processes included in the fit.
We start by discussing the fit quality. In Tab. 4.1 we report the values of the experimental $\chi^{2}$ per data point for the global fit with and without HT and PCs, and for fits with perturbative calculations at NNLO, NNLO+MHOU, and aN3LO+MHOU. Together with the global $\chi^{2}$, values are also displayed for each separate process type. In order to disentangle the effect of the shifts in the theory predictions from the effect of the additional uncertainty due to the theory covariance matrix in eq. 4.2, we performed a series of fits where only the shifts are included in the theory predictions. The corresponding $\chi^{2}$ values are also reported in Tab. 4.1. All sources of correlations are fully accounted for, and the $t_{0}$ method [Ball:2009qv] is used for multiplicative uncertainties. We refer to Ref. [NNPDF:2024nan] for further details on the inclusion of the experimental uncertainties in the $\chi^{2}$.
Inspecting Tab. 4.1, we observe a uniform improvement of the global $\chi^{2}$ upon including the HT corrections across all theory settings. Furthermore, the improvement is retained when the theory covariance matrix as in eq. 4.2 is removed, showing that the posteriors determined in section 3 are indeed moving the theory closer to the experimental data. Of course, the lowest global $\chi^{2}$ is obtained when both the shifts and the theory covariance matrix are included, as expected. A closer inspection of Tab. 4.1 reveals that at aN3LO, when the shift is applied to the inclusive jet and dijet data, the description of the top pair data also improves, indicating that the power corrections resolve some of the tension between the data for these processes. A similar effect can be seen in the NC DY data, the description of which is improved following the shift in NC DIS from the HT.
4.3 PDFs and Parton Luminosities
Figure 4.1: The PDFs with HT and PCs (in blue) for perturbative calculations at NNLO (left column), NNLO+MHOU (central column) and aN3LO+MHOU (right column), for up, anti-up, down, and anti-down. These are compared to the corresponding PDFs without HT and PCs (in red). Bands with dashed contours correspond to one-sigma uncertainties, whereas bands with continuous contours correspond to 68% confidence-level uncertainties.
Figure 4.2: Similar to fig. 4.1, but for the gluon (upper row), strange (middle row) and charm (lower row). Note the linear $x$-scale used for the strange and charm PDFs. Bands with dashed contours correspond to one-sigma uncertainties, whereas bands with continuous contours correspond to 68% confidence-level uncertainties.
We now turn to the impact of the HT and PCs on the PDFs, which are shown in fig. 4.1 for the light quarks, and in fig. 4.2 for the gluon, strange and charm. Results are displayed as functions of $x$ at $Q=1.65$ GeV, incorporating both shifts for the HT and PCs, and their (correlated) uncertainties, for the three perturbative theory settings. Overall, the impact of the HT and PCs on the PDFs is very modest, almost always within one sigma. In particular, the HT has very little impact on the large-$x$ PDFs, other than a slight increase in strangeness, and a shift in the valence charm peak to higher values of $x$. This is reassuring: the cuts to DIS data employed in NNPDF4.0 are very effective at reducing the sensitivity of the PDFs to HT. The main effect of the HT is at small $x$, where it increases PDF uncertainties due to the deweighting of the low-$x$ HERA data. This is due to the increased theoretical uncertainty from HT in the predictions for these very precise data, as was noted in Fig. 3.2. However, the increased uncertainty seen at NNLO is already reduced when MHOUs are added, and is smallest in the aN3LO fit: most of the effect seen at NNLO is not due to genuine HT, but rather to the N3LO corrections, which are substantial at small $x$.
The most interesting changes are in the gluon distribution. Again there is very little impact of the HT at large $x$, and at NNLO an increase in the uncertainty at small $x$ which is reduced once MHOUs are included, but increases again at aN3LO. More interesting is the change in the gluon at intermediate values of $x$, from roughly $0.05$ down to $0.005$, due mainly to the PCs in the single inclusive jet data (as we saw previously, the effect of PCs in the dijet data is relatively small). As remarked in the $\chi^{2}$ discussion, the PCs reduce tensions between these data and the top data. At NNLO this leads to a reduction by almost two sigma in the gluon at $x\sim 0.02$, resulting in a smoother profile, less sharply peaked. Once MHOUs are included this effect is reduced, to around one sigma, since the MHO already allows the smoother profile even before the PCs are added. At N3LO the effect of the PC is even less: the flattening of the gluon is achieved without the PC. This second observation shows that much of the tension between single inclusive jets and top at NNLO is resolved by using N3LO evolution, since both the NNLO+MHOU fit and the aN3LO+MHOU fit use the same NNLO result for the single inclusive jet partonic cross-section, with the same estimate for its uncertainty, and a similar linear PC Fig. 3.4.
Figure 4.3: Comparison of luminosities at $\sqrt{s}=13~\mathrm{TeV}$ for quark-quark, quark-antiquark, gluon-quark, and gluon-gluon channels (top to bottom) for NNLO, NNLO+MHOU, and aN3LO+MHOU (left to right), showing the ratio of the results from fits with shifts in the theory predictions due to HT and PCs (blue) to those without (red). Bands with correspond to 68% confidence-level uncertainties.
In Fig. 4.3, the luminosities in the quark-quark, quark-antiquark, gluon-gluon, and gluon-quark channels are shown at $13$ TeV, as ratios of the fits with HT and PCs to the corresponding NNLO, NNLO+MHOU, and aN3LO+MHOU [NNPDF:2024nan] standard fits. As expected from the PDF plots, the effects of HT and PCs at both high and low invariant masses are modest, well within one sigma, and largely disappear once MHOUs are added, and in the aN3LO fit. In fact the only really significant change due to the inclusion of PCs is in the gluon-gluon channel at invariant masses around the Higgs mass. This is particularly noticeable in the standard NNLO fit: here there is a one percent drop in the gluon-gluon luminosity for invariant masses between $100$ and $250$ GeV, just over two sigma. Once MHOUs are included the significance of this effect is reduced, and the aN3LO corrections reduce the impact to well within one sigma. A related effect is seen in the gluon-quark channel, though here it is never more than one sigma.
4.4 Predictions for LHC observables
Figure 4.4: Total cross-sections at $13.6$ TeV for vector boson production (upper left), production of Higgs in association with a vector boson (upper right), Higgs in gluon fusion (bottom left) and Higgs in vector boson fusion (bottom right). Each is computed with matrix element and PDFs evaluated consistently at NNLO, NNLO+MHOU and N3LO+MHOU, and with and without HT and PCs included in the determination of the PDFs. Error bars include only PDF uncertainties: in particular, the MHOU of the ME is not included.
Figure 4.5: Inclusive Higgs production cross-section in gluon fusion with the matrix element evaluated at N3LO, but with the PDFs at NNLO, NNLO+MHOU, N3LO+MHOU PDFs without HT and PCs (left), with the PDFs including the HT corrections (centre), and finally with the PDFs including PCs for single-inclusive jets and dijets (right). Error bars include only PDF uncertainties: in particular, the MHOU of the ME is not included.
We next analyse the impact of power corrections and higher twist on various totally inclusive LHC observables related to DY and Higgs production, specifically vector boson production, associated production of Higgs with a vector boson, then Higgs production in gluon fusion, and in vector boson fusion (VBF). The predictions use hard cross-sections computed up to N3LO using n3loxs [Baglio:2022wzu] for associated production, ggHiggs [Bonvini:2014jma] for gluon fusion, and proVBFH code [Dreyer:2018qbw] for VBF.
In Fig. 4.4 we collect the results of these calculations at NNLO, NNLO+MHOU, and aN3LO +MHOU, with the matrix element and PDFs computed at the same order in each case. The results with and without HT and PCs in the PDFs are then compared (no PCs are known for the matrix elements). In each case, the effect of the HT+PCs in the PDFs is small, less than one sigma. For vector boson production and Higgs production (in gluon fusion and VBF) the effect is less at N3LO than at NNLO, while for associated production the effect is a little larger at N3LO than at NNLO. In some cases there is a slight increase in uncertainty once the HT and PCs are included.
The largest effect of the HT and PCs is seen in NNLO Higgs production by gluon fusion, as expected from the discussion of the gluon-gluon luminosity in the previous section. Here, the HT+PCs reduce the total cross-section by around 1% (though still under one sigma). However when MHOUs are included, this effect drops to around 0.5%, and at N3LO it is almost negligible.
To explore this further, in Fig. 4.5 we show Higgs production in gluon fusion with the matrix element held fixed (at N3LO), but with the PDFs computed at NNLO, NNLO+MHOU, aN3LO+MHOU, then in each case with only the HT added, and then finally with both the HT and PCs included (to single inclusive jets and dijets). Without HT or PCs, going from NNLO to aN3LO PDFs reduces the cross-section by 2.1% (almost three sigma). When the HT is included, the NNLO result drops by around 0.4%, while the N3LO goes up (by 0.4%), so the difference is now only 1.3%. When the PCs are added, the NNLO result drops again, by 0.8%, while the N3LO only drops by 0.4%, so the difference between NNLO and N3LO PDFs is now only 0.9% (a little over one sigma). Therefore including HT and PCs significantly reduces the difference between using NNLO or aN3LO PDFs in combination with the N3LO matrix element.
4.5 Effect on the Strong Coupling
Figure 4.6: Results for the strong coupling $\alpha_{s}(m_{Z})$ determined in the global fit, with no HT or PC (blue), with HT but no PC (green), and with both HT and PC (red). Results are given at NNLO, NNLO+MHOU and N3LO+MHOU, in each case with and without constraints of PDF positivity.
We finally consider the effect of HT and PCs on the extraction of the strong coupling at the mass of the $Z$-boson, $\alpha_{s}(m_{Z})$, from the global dataset. We again use the theory covariance methodology, precisely as in Ref. [Ball:2025xgq], but now including also the HT and PC contributions to the theory covariance matrix. The results are shown in Fig. 4.6.
In the NNLO fit, without MHOU, the HT reduces $\alpha_{s}$ a little, while the jet PCs increase it, in each case by around one standard deviation. When MHOU is added, the best fit value of $\alpha_{s}$ drops significantly, and its uncertainty increases substantially, as found in Ref. [Ball:2025xgq]. Again the HT reduces $\alpha_{s}$, and the PCs increase it, by about one standard deviation. However, at aN3LO the pattern changes: while the aN3LO corrections reduce $\alpha_{s}$ a little, the HT correction now results in a substantial increase (more than a standard deviation). We think that this is because at aN3LO the HT is less contaminated by MHO perturbative corrections to the leading twist, and its effect on $\alpha_{s}$ is thus less easily hidden by the MHOU. The effect of the jet PCs is positive as before, but smaller (the jet cross-sections are still evaluated at NNLO+MHOU in the aN3LO fits). The result is a value of $\alpha_{s}(m_{Z})$ at aN3LO which is very consistent with that obtained at NNLO, but with substantially larger uncertainty.
As we noted in Ref. [Ball:2025xgq], all these results are subject to a further uncertainty due to the implementation of positivity constraints on the PDF fit. Since any PDF positivity constraint necessarily introduces a nongaussianity in the distribution of the PDFs, the closure testing of the methodology only works when the positivity constraints are removed. As can be seen in Fig. 4.6, releasing the positivity constraints results in lower values of $\alpha_{s}$ consistently across all the various determinations.
5 Conclusions
We have presented a determination of higher twist corrections in DIS, and linear power corrections in single-inclusive jets and dijets in a joint fit with global PDFs. We find good agreement with earlier results for higher twist. This is the first time a determination of power corrections in jets has been performed in the context of PDF fits, but we find agreement with theoretical expectations. More specifically, while the higher twist corrections are mainly taken care of by cuts in $Q^{2}$ of a few GeV2, and in $W^{2}$ at a few tens of GeV2, we find that the effect of the linear power corrections in jets can extend up to hard scales of several hundred GeV. This means in particular that when incorporating LHC data on single-inclusive jets in global PDF fits, it may be necessary for PDFs accurate to within 1% to either cut some of the low $p_{T}$ data, or include a linear power correction in the theoretical predictions (or indeed both, similar to the strategy for DIS).
We have produced PDFs which include both higher twist for DIS and linear power corrections for single-inclusive jets and dijets in the theoretical predictions used in the fit, along with their corresponding uncertainties encoded in a theory covariance matrix. We have assessed the impact of these corrections on a series of representative LHC observables and found that while they are for the most part small, they can be as large as one percent for processes sensitive to the gluon-gluon luminosity in the intermediate mass region, such as Higgs in gluon fusion. We have likewise examined the effect on the value of $\alpha_{s}$ determined from hadronic processes, and again found effects as large as one percent.
Since the effects of higher twist and linear power corrections in jets can be relevant at the current level of precision for various processes, they should be accounted for in future global PDF determinations that hope to achieve a precision of 1% in LHC observables. The empirical determination of power corrections in Drell-Yan and vector boson production processes, and the possible linear power corrections in top production [Makarov:2023uet], are left for future studies.
The PDF sets with power corrections are made available in LHAPDF format [Buckley:2014ana] on the NNPDF website
https://nnpdf.mi.infn.it/nnpdf4-0-higher-twist/
under the following names:
NNPDF40_nnlo_as_01180_ht
NNPDF40_nnlo_as_01180_mhou_ht
NNPDF40_an3lo_as_01180_mhou_ht .
After this work was completed, two other studies of higher twist in the context of PDF fits and $\alpha_{s}$ appeared [Harland-Lang:2025wvm, Alekhin:2025qdj]. Neither of these consider higher twist for charge current processes, or linear power corrections to jet processes.
Acknowledgements:
We would like to thank Gavin Salam for discussions at Gargnano in September 2023, where he emphasised the possible significance of linear power corrections to jet observables in global PDF fits. We would also like to thank the members of the NNPDF collaboration for insightful discussions during the course of this work. Finally, we are grateful to Felix Hekhorn for providing valuable feedback on the first draft of this manuscript. R.D.B. and R.S. are supported by the Science and Technology Facilities Council (STFC) via grant awards ST/T000600/1 and ST/X000494/1.

### v2

@@
 We determine these corrections and their associated correlated uncertainties using a methodology based on the theory covariance formalism, previously used to account for nuclear uncertainties and missing higher order uncertainties (MHOUs) in global PDF determinations.
-We then study the impact of the power corrections on the extracted PDFs, and demonstrate an improved description of the datam due to a reduced sensitivity to DIS data in the low-$x$ region where higher twist uncertainties are relatively large, and a reduced sensitivity to single inclusive jet data at relatively low $p_{T}$, where linear power corrections can be significant.
+We then study the impact of the power corrections on the extracted PDFs, and demonstrate an improved description of the data due to a reduced sensitivity to DIS data in the low-$x$ region where higher twist uncertainties are relatively large, and a reduced sensitivity to single inclusive jet data at relatively low $p_{T}$, where hadronization corrections can be large.
 Finally, we assess the impact of power corrections on observables relevant to LHC phenomenology, including Higgs production via gluon fusion, and the determination of $\alpha_{s}$.
@@
 Nevertheless, some significant sources of theoretical uncertainty remain, both perturbative and non-perturbative, and in both the hard cross-sections themselves and in the parton distribution functions (PDFs).
-If our theoretical predictions are to match the accuracy and precision of LHC data, we need to control all theoretical uncertainties at the one percent level [Amoroso:2022eow].
+If our theoretical predictions are to match the accuracy and precision of LHC data, we need to control all theoretical uncertainties at the one percent level [16].
 A systematic estimation of all possible sources of significant theoretical uncertainty is important for two main reasons.
@@
 Second, and more subtly, since theoretical predictions at hadron colliders rely on accurate and precise PDFs, it is necessary to also account systematically for all the various theoretical uncertainties in the calculations used in global PDF fits.
-Theoretical uncertainties arising from missing higher-orders (MHOUs) have been included in PDF fits using a theory covariance matrix [Ball:2018lag] constructed from scale variations [NNPDF:2019vjt, NNPDF:2019ubu, NNPDF:2024dpb], or by fitting nuisance parameters [McGowan:2022nag].
-The covariance matrix approach has also been used to incorporate the non-perturbative uncertainties due to nuclear effects in global fits with deep-inelastic scattering (DIS) and Drell-Yan (DY) data on heavy nuclei and deuteron targets [Ball:2018twp, Ball:2020xqw].
+Theoretical uncertainties arising from missing higher-orders (MHOUs) have been included in PDF fits using a theory covariance matrix [29] constructed from scale variations [5, 7, 35], or by fitting nuisance parameters [69].
+The covariance matrix approach has also been used to incorporate the non-perturbative uncertainties due to nuclear effects in global fits with deep-inelastic scattering (DIS) and Drell-Yan (DY) data on heavy nuclei and deuteron targets [30, 31].
 Besides theoretical uncertainties due to perturbative corrections and nuclear effects, there are further uncertainties due to effects which are suppressed by a power of the hard scale.
@@
 Additional terms, known as higher twist (HT) corrections, are always present, but suppressed by powers of the hard scale $Q^{2}$.
-These HT corrections are in essence non-perturbative, and over the years there have been many attempts to determine them empirically by fitting $\mathcal{O}(\Lambda^{2}/Q^{2})$ corrections (here $\Lambda$ is the QCD scale) in the context of a PDF fit [Virchaux:1991jc, Alekhin:1998df, Alekhin:1999af, Martin:1998np, Martin:2003sk, Blumlein:2008kz, Alekhin:2012ig, Ball:2013gsa, Thorne:2014toa, Alekhin:2017kpj, Alekhin:2022tip, Cerutti:2025yji].
+These HT corrections are in essence non-perturbative, and over the years there have been many attempts to determine them empirically by fitting $\mathcal{O}(\Lambda^{2}/Q^{2})$ corrections (here $\Lambda$ is the QCD scale) in the context of a PDF fit [75, 15, 13, 67, 66, 43, 11, 24, 73, 10, 14, 51, 63, 12].
 This is challenging, as the HT corrections can easily be confused with the unknown terms behaving as powers of $\mathcal{O}\left(1/\log(Q^{2}/\Lambda^{2})\right)$ arising due to MHOUs.
 Consequently most global PDF fits impose kinematic cuts on DIS data to ensure that the HT corrections are small.
-For hadronic processes, such as DY or inclusive jets, there is no operator product expansion, but factorization still holds in the limit of large momentum transfer [Collins:1989gx].
+For hadronic processes, such as DY or inclusive jets, there is no operator product expansion, but factorization still holds in the limit of large momentum transfer [54].
 Again there are non-perturbative contributions suppressed by powers of the hard scale $Q_{T}$.
-For sufficiently inclusive processes with a color singlet final state, such as DY, the power corrections (PCs) are expected to be $\mathcal{O}(\Lambda^{2}/Q_{T}^{2})$ [Beneke:1995pq, Caola:2021kzt], and thus small except close to threshold.
+For sufficiently inclusive processes with a color singlet final state, such as DY, the power corrections (PCs) are expected to be $\mathcal{O}(\Lambda^{2}/Q_{T}^{2})$ [41, 49], and thus small except close to threshold.
 However for many other hadronic observables there are generally linear power corrections of $\mathcal{O}(\Lambda/Q_{T})$.
-For example, linear PCs to jet processes have been studied extensively in $e^{+}e^{-}$ annihilation [Dasgupta:2003iq, Dasgupta:2007wa, Dasgupta:2008di, Cacciari:2007aq], with analytic estimates of the shift in the jet energy due to hadronization and underlying event.
-Studies of jets in hadronic collisions [Olness:2009qd] have extended these analyses to estimate the size of theoretical uncertainties related to various non-perturbative effects in inclusive single jet cross-sections.
+For example, linear PCs to jet processes have been studied extensively in $e^{+}e^{-}$ annihilation [58, 56, 57, 46], with analytic estimates of the shift in the jet energy due to hadronization and underlying event.
+Studies of jets in hadronic collisions [70] have extended these analyses to estimate the size of theoretical uncertainties related to various non-perturbative effects in inclusive single jet cross-sections.
 These estimates suggest that the linear PCs to jet processes might be significant even up to scales as large as $100$ GeV.
-Linear PCs are also expected in top production [Makarov:2023uet], unless one uses a short distance scheme for the top quark mass.
+Linear PCs are also expected in top production [65], unless one uses a short distance scheme for the top quark mass.
 Since linear PCs fall off much more slowly with hard scale than HT, it is even more difficult to disentangle them from MHOUs.
 To the best of our knowledge, an empirical determination of linear PCs to jets in the context of a global PDF fit has not yet been attempted.
-In this paper, we use the theory covariance matrix formalism developed in [Ball:2018lag, Ball:2021icz], recently validated in Ref.
-[Ball:2025xgq], to provide reliable estimates of the corrections and theoretical uncertainties due to HT in DIS, and to PCs in single jet inclusive cross-sections and dijet cross-sections.
-The analysis is performed entirely within the open-source NNDPF framework [NNPDF:2021uiq].
+In this paper, we use the theory covariance matrix formalism developed in [29, 38], recently validated in Ref.
+[20], to provide reliable estimates of the corrections and theoretical uncertainties due to HT in DIS, and to PCs in single jet inclusive cross-sections and dijet cross-sections.
+The analysis is performed entirely within the open-source NNDPF framework [33].
 The paper is organised as follows.
@@
 In section 4, we will include these corrections and uncertainties in a global PDF fit, assess its impact on the PDFs and relative luminosities, on the theoretical predictions for fundamental LHC processes such as the Higgs boson production, and on the determination of the strong coupling $\alpha_{s}$ from hadronic data.
-Finally, in section 5 we summarise our main conclusions and provide information on how to access the PDF fits with HT and PCs in LHAPDF format [Buckley:2014ana].
+Finally, in section 5 we summarise our main conclusions and provide information on how to access the PDF fits with HT and PCs in LHAPDF format [45].
 2 Methodology for the Determination of Power Corrections
@@
 This is based on the general formalism developed in Ref.
-[Ball:2021icz], where it is shown how to account for correlations between theory uncertainties in the PDFs and in the matrix element used in a prediction.
+[38], where it is shown how to account for correlations between theory uncertainties in the PDFs and in the matrix element used in a prediction.
 When applied to parameters of physical interest, the methodology allows for the extraction of such parameters, together with their correlated uncertainties, in such a way as to account for all the correlations with the underlying PDFs.
-As such, the methodology was recently validated in a determination of the strong coupling [Ball:2025xgq], by comparing results to those obtained using the more cumbersome correlated replica method [Ball:2018iqk], and by closure testing [NNPDF:2017mvq].
+As such, the methodology was recently validated in a determination of the strong coupling [20], by comparing results to those obtained using the more cumbersome correlated replica method [26], and by closure testing [32].
 In the following, we provide a brief overview of the theory covariance methodology before moving to its specific application to power corrections.
@@
 Theoretical uncertainties and their correlations can be incorporated in the fitting procedure through the implementation of a theoretical covariance matrix $C^{\textrm{th}}$.
-If we assume that these uncertainties are Gaussian and independent of the experimental uncertainties, the total covariance matrix $C$ can be written as the sum of the experimental $t_{0}$ covariance matrix [Ball:2009qv] $C^{\textrm{exp}}_{t_{0}}$ and a theory covariance matrix [Ball:2018lag, Ball:2018twp, Ball:2020xqw, NNPDF:2019vjt, NNPDF:2019ubu, NNPDF:2024dpb] incorporating in particular all MHOUs and nuclear uncertainties
+If we assume that these uncertainties are Gaussian and independent of the experimental uncertainties, the total covariance matrix $C$ can be written as the sum of the experimental $t_{0}$ covariance matrix [28] $C^{\textrm{exp}}_{t_{0}}$ and a theory covariance matrix [29, 30, 31, 5, 7, 35] incorporating in particular all MHOUs and nuclear uncertainties
 $C_{ij}=C^{\textrm{exp}}_{t_{0},ij}+C^{\textrm{th}}_{ij}\,$, (2.1)
@@
 The figure of merit in Eq.
-(2.3) ensures that the relative weights of the data points during the fit account for both experimental and theoretical uncertainties [Ball:2018lag].
+(2.3) ensures that the relative weights of the data points during the fit account for both experimental and theoretical uncertainties [29].
 The $\chi^{2}$ in Eq.
@@
 Following Ref.
-[Ball:2021icz], additional sources of theoretical uncertainty can be modelled as fully correlated shifts in the predictions induced by a set of $N_{\textrm{nuis}}$ sources of theoretical uncertainty
+[38], additional sources of theoretical uncertainty can be modelled as fully correlated shifts in the predictions induced by a set of $N_{\textrm{nuis}}$ sources of theoretical uncertainty
 $T({\lambda})=T+\sum_{\alpha=1}^{N_{\textrm{nuis}}}\lambda_{\alpha}\beta_{\alpha}\,$, (2.5)
@@
 (2.15) are linear in the nuisance parameters, there is no need (unlike in Refs.
-[AbdulKhalek:2019bux, AbdulKhalek:2019ihb, NNPDF:2024dpb, Ball:2025xgq]) to consider contributions from positive and negative shifts separately, nor to consider complicated prescriptions: Eq.
+[4, 6, 35, 20]) to consider contributions from positive and negative shifts separately, nor to consider complicated prescriptions: Eq.
 (2.7) is perfectly sufficient.
@@
 We do not include DIS data on charm or bottom production in this study, nor data on DIS+jets.
-There are two distinct approaches to modelling higher twist corrections to theoretical predictions for DIS data: the additive approach [Alekhin:1998df, Alekhin:1999af, Blumlein:2008kz, Alekhin:2012ig, Alekhin:2017kpj, Alekhin:2022tip], and the multiplicative approach [Virchaux:1991jc, Martin:1998np, Martin:2003sk, Ball:2013gsa, Thorne:2014toa].
+There are two distinct approaches to modelling higher twist corrections to theoretical predictions for DIS data: the additive approach [15, 13, 43, 11, 10, 14], and the multiplicative approach [75, 67, 66, 24, 73].
 In the additive approach one ignores all perturbative corrections to the HT.
@@
 Both approaches are thus approximate, and which one adopts is largely a matter of taste (see Ref.
-[Cerutti:2025yji] for a recent phenomenological comparison).
+[51] for a recent phenomenological comparison).
 We consider HT corrections as multiplicative shifts to DIS observables, which comprise measurements of structure functions $F_{2}$ and reduced cross-sections.
 We model the shifts as functions of Bjorken-$x$, to be determined empirically.
-Moreover, given that in principle HT can depend on nuclear effects, the shifts for proton and deuteron targets are parametrised independently [Cerutti:2025yji].
+Moreover, given that in principle HT can depend on nuclear effects, the shifts for proton and deuteron targets are parametrised independently [51].
 Hence, we write the measured structure functions as
@@
 $\sigma_{\rm NC}(x,Q^{2})=\sigma^{\textrm{LT}}_{\rm NC}(x,Q^{2})\left(1+\frac{H_{2}^{p}(x)}{Q^{2}}\right)\,.$ (2.20)
-For ratio data, we have
+Here we make the simplifying assumption that higher twist corrections to $F_{L}$ is the same as for $F_{2}$, since the data are not good enough to determine them independently. For ratio data, we have
 $\frac{F_{2}^{d}}{F_{2}^{p}}=\left(\frac{F_{2}^{d}}{F_{2}^{p}}\right)^{LT}\left(1+\frac{H_{2}^{(d)}-H_{2}^{(p)}}{Q^{2}+H_{2}^{(p)}}\right)\,.$ (2.21)
@@
 Finally, DIS CC cross-sections are corrected similarly, by introducing a common multiplicative shift for CHORUS, NuTeV, and HERA CC data regardless of the nuclear target used in the experiment.
-Thus we will assume that nuclear effects at higher twist in CC data are similar to those included at leading twist [Ball:2018twp].
-This approximation will be justified a priori.
+Thus we will assume that CC $F_{2}$, $F_{3}$ and $F_{L}$ are all corrected by the same factor, and furthermore that nuclear effects at higher twist in CC data are similar to those included at leading twist [30].
 Hence, the shift for any CC cross-section will be written as
@@
 Here $f_{p}(x_{1})$ and $f_{q}(x_{2})$ are the PDFs of the incoming partons $p$ and $q$, and $d\hat{\sigma}_{pq\rightarrow\textrm{jet}}/dp_{T}dy$ is the partonic cross-section computed at NNLO in perturbative QCD.
-MHOUs estimated from scale variations have been accounted for in a variant of NNPDF4.0 fits [NNPDF:2024dpb].
+MHOUs estimated from scale variations have been accounted for in a variant of NNPDF4.0 fits [35].
 While it has been understood for some time that there are linear power corrections to these processes, from a wide variety of sources, in global PDF fits these are generally ignored.
@@
 As a results, the measured energy of the jet is shifted due to the external contribution from the underlying event that flows into the jet.
-Investigations to disentangle the underlying event contribution from single-inclusive jets exploiting Monte Carlo event generators have been performed by ATLAS [ATLAS:2014iez] and CMS [CMS:2013ycn].
+Investigations to disentangle the underlying event contribution from single-inclusive jets exploiting Monte Carlo event generators have been performed by ATLAS [3] and CMS [52].
 On the other hand, in the process of hadronization, some of the jet energy leaks out from the jet cone, reducing the energy that is eventually measured.
 This process is modelled through the use of final state Monte Carlo generators.
-Experimentalists correct for these effects in their measurements, with associated systematic uncertainties.
+Experimentalists can correct for these effects in their measurements, by computing a correction factor with associated systematic uncertainties.
 However in addition to all this, the theoretical calculation of the perturbative jet cross-section is also subject to non-perturbative uncertainties, resulting in power corrections of $O(\Lambda/p_{T})$ to the parton level transverse momentum and thus to the perturbative cross-sections.
-These genuine non-perturbative effects are by their nature very difficult to disentangle from the quasi-partonic effects already corrected for in the experimental measurements and thus, as with higher twist, an empirical approach is the only practical way to assess them.
+These genuine non-perturbative effects are by their nature very difficult to disentangle from the quasi-partonic effects estimated by using final state Monte Carlos, and thus, as with higher twist, an empirical approach is the most practical way to assess them quantitatively.
 Whatever the source of the effect, a change in energy results in a shift $\delta p_{T}$ to the fixed parton level transverse momentum $p_{T}$ of the jet.
 The correction $\delta p_{T}$ has been estimated in Refs.
-[Dasgupta:2007wa, Dasgupta:2008di, Cacciari:2007aq].
-Since jet cross-sections fall very steeply with $p_{T}$ (typically as $\sim p_{T}^{-n}$, with $n$ as large as six or seven), such a shift results in a multiplicative power correction of the form $1+n\delta p_{T}/p_{T}$, which may be large simply because $n$ is large [Olness:2009qd].
-We will thus model the shift in the single-inclusive jet cross-section induced by power corrections as
+[56, 57, 46].
+Since jet cross-sections fall very steeply with $p_{T}$ (typically as $\sim p_{T}^{-n}$, with $n$ as large as six or seven), such a shift results in a multiplicative power correction of the form $1+n\delta p_{T}/p_{T}$, which may be large simply because $n$ is large [70].
+We will thus model the shift in the single-inclusive jet cross-section induced by all these various power corrections as
 $\frac{d\sigma_{\textrm{jet}}}{dp_{T}dy}=\left[\frac{d\sigma_{\textrm{jet}}}{dp_{T}dy}\right]_{\text{pert}}\biggl(1+\frac{H_{j}(y)}{p_{T}}\biggr)\,$, (2.25)
@@
 This problem is compounded by the fact that the N3LO corrections to jet processes are unknown, to the perturbative predictions are less precise than for DIS.
-On the other hand, since as explained in [Olness:2009qd] the scale of the linear power corrections may be relatively large (several GeV rather than several hundred MeV), they may continue to be significant up to relatively high scales.
+On the other hand, since as explained in [70] the scale of the linear power corrections may be relatively large (several GeV rather than several hundred MeV), they may continue to be significant up to relatively high scales.
 They can thus be significant, even at the LHC.
@@
 All these results will be obtained using the theory covariance methodology described in the previous section.
-The fits presented in this section and the next have been produced using the open-source NNPDF framework [NNPDF:2021uiq] based on the same dataset and theory parameters as used in NNPDF4.0 [NNPDF:2021njg] with MHOUs as in Ref.
-[NNPDF:2024dpb] and aN3LO corrections as in Ref.
-[NNPDF:2024nan].
+The fits presented in this section and the next have been produced using the open-source NNPDF framework [33] based on the same dataset and theory parameters as used in NNPDF4.0 [34] with MHOUs as in Ref.
+[35] and aN3LO corrections as in Ref.
+[37].
 Theory predictions have been generated with the pipeline of Ref.
-[Barontini:2023vmr], which supersedes the tools used for NNPDF4.0.
-In particular, interpolation grids are generated using PineAPPL [Carrazza:2020gss] with the DGLAP evolution computed using EKO [Candido:2022tld] and DIS observables using the Yadism package [Candido:2024rkr].
+[39], which supersedes the tools used for NNPDF4.0.
+In particular, interpolation grids are generated using PineAPPL [50] with the DGLAP evolution computed using EKO [48] and DIS observables using the Yadism package [47].
 For a detailed discusion of the updated theory pipeline see App A of Ref.
-[NNPDF:2024djq].
+[36].
 3.1 Determination of Higher Twist
-The DIS measurements included in NNPDF4.0 [NNPDF:2021njg] consist of fixed-target NC structure function data from SLAC [Whitlow:1991uw], BCDMS [Benvenuti:1989rh], and NMC [Arneodo:1996kd, Arneodo:1996qe], fixed-target inclusive and dimuon CC cross-sections from CHORUS [Onengut:2005kv] and NuTeV [Goncharov:2001qe, MasonPhD], and collider NC and CC cross-sections from HERA [Abramowicz:2015mha].
+The DIS measurements included in NNPDF4.0 [34] consist of fixed-target NC structure function data from SLAC [76], BCDMS [42], and NMC [17, 18], fixed-target inclusive and dimuon CC cross-sections from CHORUS [71] and NuTeV [62, 68], and collider NC and CC cross-sections from HERA [9].
 In NNPDF4.0 (and indeed all earlier NNPDF global fits) cuts are applied to all DIS data, $Q^{2}>3.49~{\rm GeV}^{2}$, $W^{2}>12.5~{\rm GeV}^{2}$, in order to avoid the kinematic regions where the HT is expected to be significant.
-All calculations of DIS processes in NNPDF4.0 include target mass corrections [Schienbein:2007gr], following the implementation of Refs.
-[Georgi:1976ve, Ball:2008by], heavy quark mass corrections implemented using FONLL [Forte:2010ta, Ball:2011mu, Barontini:2024xgu], fitted charm [Ball:2015tna, Ball:2015dpa, Ball:2016neh], and nuclear corrections and uncertainties, both for heavy nuclei [Ball:2018twp] and deuteron targets [Ball:2020xqw].
+All calculations of DIS processes in NNPDF4.0 include target mass corrections [72], following the implementation of Refs.
+[61, 27], heavy quark mass corrections implemented using FONLL [60, 23, 40], fitted charm [22, 25, 21], and nuclear corrections and uncertainties, both for heavy nuclei [30] and deuteron targets [31].
 We estimate the higher twist using the theory covariance methodology, as described in section 2, constructing a prior theory covariance matrix by computing the shifts in DIS predictions for a prior variation of the HT parameters $h_{\alpha}$.
@@
 In particular, at $x=0.9$ there is no constraint at all, with the posterior equal to the prior.
-This said, we see no evidence here for the expectation [Dasgupta:1996hh], based on renormalons, that HT effects for CC processes are larger than for NC processes, or that they have a different shape.
+This said, we see no evidence here for the expectation [55], based on renormalons, that HT effects for CC processes are larger than for NC processes, or that they have a different shape (although it should be remembered that in our analysis sum rule constraints have been imposed only at leading twist).
 In Fig.
-3.3 we show a comparison of the HT for the $F_{2}$ structure function in proton and deuteron determined in our global NNLO fit, with results obtained in an early PDF fit to just SLAC and BCDMS data by Virchaux and Milsztajn [Virchaux:1991xp].
+3.3 we show a comparison of the HT for the $F_{2}$ structure function in proton and deuteron determined in our global NNLO fit, with results obtained in an early PDF fit to just SLAC and BCDMS data by Virchaux and Milsztajn [74].
 While this fit was only at NLO, it used a similar multiplicative parametrization to our own, and target mass corrections were included, so the comparison is meaningful.
 Considering the fact that this fit was performed before the advent of NMC and HERA data, and all the technical improvements made over the intermediate thirty years, the agreement is rather impressive, though it should not come as too much of a surprise since it is the same data (SLAC and BCDMS) that dominates the large-$x$, low-$Q^{2}$ region today.
-Direct comparisons with more recent determinations which use an additive HT parametrization [Alekhin:1998df, Alekhin:1999af, Blumlein:2008kz, Alekhin:2012ig, Alekhin:2017kpj, Alekhin:2022tip] are more difficult, since the results necessarily depend to some extent on the choice of scale.
+Direct comparisons with more recent determinations which use an additive HT parametrization [15, 13, 43, 11, 10, 14] are more difficult, since the results necessarily depend to some extent on the choice of scale.
 However for these too, the results for HT correction to $F_{2}^{p}$ are qualitatively consistent with our own, within uncertainties.
-Figure 3.3: Comparison of our NNLO $F_{2}$ higher twist to that determined in protons and deuteron by Virchaux and Milsztajn in a NLO PDF fit to SLAC and BCDMS data [Virchaux:1991xp].
+Figure 3.3: Comparison of our NNLO $F_{2}$ higher twist to that determined in protons and deuteron by Virchaux and Milsztajn in a NLO PDF fit to SLAC and BCDMS data [74].
 They include TMCs, and use the same multiplicative parametrization of the HT, fitting proton and deuteron data separately, so the results are directly comparable to our own.
@@
 This shows the difficulty in distinguishing phenomenologically between $Q^{2}$ dependence which is power like, and $Q^{2}$ dependence which is merely a power of a logarithm.
-This effect was also seen when going from NLO to NNLO in earlier attempts to determine HT [Martin:2003sk].
+This effect was also seen when going from NLO to NNLO in earlier attempts to determine HT [66].
 One is forced to conclude that even the HT determined in the aN3LO fit may still be contaminated by a perturbative contribution.
-Similarly, $H_{2}^{d}$ may be tangled with the deuteron nuclear correction [Ball:2020xqw], and $H_{CC}$ with the nuclear corrections applied to much of the CC data [Ball:2018twp].
+Similarly, $H_{2}^{d}$ may be tangled with the deuteron nuclear correction [31], and $H_{CC}$ with the nuclear corrections applied to much of the CC data [30].
 The size and significance of the HT we determine is also displayed in the kinematic plots in Fig.
@@
 We now turn to the determination of power corrections for jet observables.
-The jet cross-sections included in the NNPDF4.0 fit are double differential single-inclusive jet data at $\sqrt{s}=8$ TeV and dijet cross-sections at $\sqrt{s}=7$ TeV from ATLAS [ATLAS:2017kux, ATLAS:2013jmu] and CMS [CMS:2016lna, CMS:2012ftr].
-No kinematic cuts were applied to these data in NNPDF4.0.
+The jet cross-sections included in the NNPDF4.0 fit are double differential single-inclusive jet data at $\sqrt{s}=8$ TeV and dijet cross-sections at $\sqrt{s}=7$ TeV from ATLAS [1, 2] and CMS [64, 53], following a detailed study of the reliability and impact of the datasets [8].
+The jets are constructed in both cases using the anti-$k_{T}$ algorithm, and the data used in the fit were those with $R=0.6$ for ATLAS, $R=0.7$ for CMS.
+The published datasets are at hadron level.
+Hadronization, underlying event and multi-parton corrections were estimated by the experimentalists from the spread in several runs of final state Monte Carlos with different tunes and input PDFs, was included in the covariance matrix.
+However the spread in the ATLAS estimates is very wide (consistent with zero), while that from CMS is much narrower, and significantly nonzero, particularly at low scale.
+Given this fact, it was decided in NNPDF4.0 not to correct the central values of the data, but to implement only an overall hadronization uncertainty as a contribution to the experimental covariance matrix.
 An internal cut on the jet $p_{T}$ was applied to the single-jet measurements of CMS in order to avoid points for which $p_{T}\lesssim 70$ GeV, where fixed-order calculations are not reliable.
 This cut is not necessary for ATLAS single-jet data since the lowest $p_{T}$ bin starts at $70$ GeV.
-Finally, no additional cuts are applied to dijet data, as the lowest value of the hard scale is one order of magnitude larger than that for the single-inclusive jet data.
-These considerations will become relevant in the discussion below.
+No additional cuts are applied to dijet data, as the lowest value of the hard scale is one order of magnitude larger than that for the single-inclusive jet data.
+All these considerations will become relevant in the discussion below.
 As discussed in Sec.
@@
 It should be noted that the previous study in Ref.
-[Olness:2009qd] estimated a positive power correction of about $7$ GeV for single-inclusive jet cross-sections, independent of rapidity.
+[70] estimated a positive power correction of about $7$ GeV for single-inclusive jet cross-sections, independent of rapidity.
 Our result is consistent with this estimate.
-For the dijets, we find only weak evidence for a linear PC in the NNLO fit, although uncertainties remain large, showing that the data provide only weak constraints on the PCs.
+For the dijets, we find rather weaker evidence for a linear PC at central rapidity in the NNLO fit, with uncertainties remaining large.
 Note that for the dijet kinematics the bins with lowest invariant mass (and thus most likely to be affected by PCs) are at central rapidity, and even here the lowest invariant mass bin is still at a rather higher scale than the lowest $p_{T}$ scale of the single inclusive jet data.
@@
 The colour code represents the ratio between the posterior uncertainty due to PCs and the experimental uncertainty.
+Figure 3.6: The hadronization correction as a function of $p_{T}$ for the 88 TeV single inclusive jet data, for $0<|y|<0.5$ (left) and $2.5<|y|<3.0$ (right).
+Our empirical estimates for the linear power correction at NNLO, NNLO+MHOU and N3LO+MHOU are compared to the ATLAS and CMS estimates obtained using final state Monte Carlo generators [1, 64].
 There is perhaps a small reduction in the size of the PC to single-inclusive jets when MHOUs are added at NNLO, particularly at large rapidities, but in the aN3LO fit the result is entirely consistent with NNLO.
@@
 The underlying reason for this is kinematic: it is easier to identify unambiguously the hard scale for dijets than for single-inclusive jets.
+Finally, it is interesting to compare these empirical estimates of linear power corrections Eq.
+2.25 with the estimates of hadronization, underlying event and multi-parton corrections made by ATLAS and CMS using final state Monte Carlos, which make no assumption regarding the $p_{T}$ dependence of the correction.
+The comparison for the CMS $8$ TeV single inclusive jet data [1, 64], for both central rapidity and high rapidity bins, is shown in Fig.
+3.6.
+At central rapidity the agreement is very good, though at low $p_{T}$ the Monte Carlo estimates rise more steeply than our own, while at high $p_{T}$ they fall less quickly.
+At high rapidity the agreement is less good, the Monte Carlo estimate being rather higher, though still consistent within uncertainties.
+The consistency between these two very different approaches is quite encouraging, suggesting that if the Monte Carlo estimates of the hadronization (and multi-particle interactions) had been used to correct the data, converting it from hadron level to parton level, the linear power corrections found in our analysis would have been rather smaller, though more dependent on rapidity.
+Similar remarks apply to the dijet datasets.
 4 The NNPDF4.0HT Parton Sets
-We evaluate the phenomenological consequences of including the HT corrections and PCs discussed in the previous sections at the level of a new global PDF fit, and discuss the impact on relevant LHC observables.
+We evaluate the phenomenological consequences of including the HT corrections and PCs discussed in the previous sections at the level of a new global PDF fit, and discuss the impact on relevant LHC observables and on the determination of $\alpha_{s}$.
 4.1 Global PDF fits with Power Corrections
@@
 However, we do not suggest performing a fit with these lowered cuts to determine the PDFs, since such a fit would include data which is significantly affected by HT.
-Instead, we use the HT and the PCs determined above to give corrections to the theory predictions in a fit with the conventional cuts used in NNPDF4.0 [NNPDF:2021njg], i.e.
+Instead, we use the HT and the PCs determined above to give corrections to the theory predictions in a fit with the conventional cuts used in NNPDF4.0 [34], i.e.
 $Q^{2}>3.49$ GeV2 and $W^{2}>12.5$ GeV2.
@@
 4.1.
-All sources of correlations are fully accounted for, and the $t_{0}$ method [Ball:2009qv] is used for multiplicative uncertainties.
+All sources of correlations are fully accounted for, and the $t_{0}$ method [28] is used for multiplicative uncertainties.
 We refer to Ref.
-[NNPDF:2024nan] for further details on the inclusion of the experimental uncertainties in the $\chi^{2}$.
+[37] for further details on the inclusion of the experimental uncertainties in the $\chi^{2}$.
 Inspecting Tab.
@@
 In Fig.
-4.3, the luminosities in the quark-quark, quark-antiquark, gluon-gluon, and gluon-quark channels are shown at $13$ TeV, as ratios of the fits with HT and PCs to the corresponding NNLO, NNLO+MHOU, and aN3LO+MHOU [NNPDF:2024nan] standard fits.
+4.3, the luminosities in the quark-quark, quark-antiquark, gluon-gluon, and gluon-quark channels are shown at $13$ TeV, as ratios of the fits with HT and PCs to the corresponding NNLO, NNLO+MHOU, and aN3LO+MHOU [37] standard fits.
 As expected from the PDF plots, the effects of HT and PCs at both high and low invariant masses are modest, well within one sigma, and largely disappear once MHOUs are added, and in the aN3LO fit.
@@
 We next analyse the impact of power corrections and higher twist on various totally inclusive LHC observables related to DY and Higgs production, specifically vector boson production, associated production of Higgs with a vector boson, then Higgs production in gluon fusion, and in vector boson fusion (VBF).
-The predictions use hard cross-sections computed up to N3LO using n3loxs [Baglio:2022wzu] for associated production, ggHiggs [Bonvini:2014jma] for gluon fusion, and proVBFH code [Dreyer:2018qbw] for VBF.
+The predictions use hard cross-sections computed up to N3LO using n3loxs [19] for associated production, ggHiggs [44] for gluon fusion, and proVBFH code [59] for VBF.
 In Fig.
@@
 We again use the theory covariance methodology, precisely as in Ref.
-[Ball:2025xgq], but now including also the HT and PC contributions to the theory covariance matrix.
+[20], but now including also the HT and PC contributions to the theory covariance matrix.
 The results are shown in Fig.
@@
 When MHOU is added, the best fit value of $\alpha_{s}$ drops significantly, and its uncertainty increases substantially, as found in Ref.
-[Ball:2025xgq].
+[20].
 Again the HT reduces $\alpha_{s}$, and the PCs increase it, by about one standard deviation.
@@
 As we noted in Ref.
-[Ball:2025xgq], all these results are subject to a further uncertainty due to the implementation of positivity constraints on the PDF fit.
+[20], all these results are subject to a further uncertainty due to the implementation of positivity constraints on the PDF fit.
 Since any PDF positivity constraint necessarily introduces a nongaussianity in the distribution of the PDFs, the closure testing of the methodology only works when the positivity constraints are removed.
@@
 We find good agreement with earlier results for higher twist.
-This is the first time a determination of power corrections in jets has been performed in the context of PDF fits, but we find agreement with theoretical expectations.
-More specifically, while the higher twist corrections are mainly taken care of by cuts in $Q^{2}$ of a few GeV2, and in $W^{2}$ at a few tens of GeV2, we find that the effect of the linear power corrections in jets can extend up to hard scales of several hundred GeV.
+This is the first time an empirical determination of power corrections in jets has been performed in the context of PDF fits, but we find agreement with theoretical expectations, and in particular with estimates of hadronization corrections made using final state Monte Carlos.
+While the higher twist corrections are mainly taken care of by cuts in $Q^{2}$ of a few GeV2, and in $W^{2}$ at a few tens of GeV2, we find that the effect of the linear power corrections in jets can extend up to hard scales of several hundred GeV.
 This means in particular that when incorporating LHC data on single-inclusive jets in global PDF fits, it may be necessary for PDFs accurate to within 1% to either cut some of the low $p_{T}$ data, or include a linear power correction in the theoretical predictions (or indeed both, similar to the strategy for DIS).
+The size of the power correction could be reduced by first correcting the data (from hadron to parton level) using a correction factor estimate using final state Monte Carlos, if such factors were determined reliably and consistently across all available jet datasets.
 We have produced PDFs which include both higher twist for DIS and linear power corrections for single-inclusive jets and dijets in the theoretical predictions used in the fit, along with their corresponding uncertainties encoded in a theory covariance matrix.
@@
 Since the effects of higher twist and linear power corrections in jets can be relevant at the current level of precision for various processes, they should be accounted for in future global PDF determinations that hope to achieve a precision of 1% in LHC observables.
-The empirical determination of power corrections in Drell-Yan and vector boson production processes, and the possible linear power corrections in top production [Makarov:2023uet], are left for future studies.
-The PDF sets with power corrections are made available in LHAPDF format [Buckley:2014ana] on the NNPDF website
+The empirical determination of power corrections in Drell-Yan and vector boson production processes, and the possible linear power corrections in top production [65], are left for future studies.
+The PDF sets with power corrections are made available in LHAPDF format [45] on the NNPDF website
 https://nnpdf.mi.infn.it/nnpdf4-0-higher-twist/
@@
 NNPDF40_an3lo_as_01180_mhou_ht .
-After this work was completed, two other studies of higher twist in the context of PDF fits and $\alpha_{s}$ appeared [Harland-Lang:2025wvm, Alekhin:2025qdj].
-Neither of these consider higher twist for charge current processes, or linear power corrections to jet processes.
 Acknowledgements:
@@
 We would also like to thank the members of the NNPDF collaboration for insightful discussions during the course of this work.
-Finally, we are grateful to Felix Hekhorn for providing valuable feedback on the first draft of this manuscript.
+Finally, we are grateful to Felix Hekhorn and Emanuele Nocera for providing valuable feedback on the earlier drafts of this manuscript.
 R.D.B.